OpenStudy (anonymous):

solving multi-step equations: -6(-5+3m)=-78 step by step pleaseeee

OpenStudy (diyadiya):

(-6*-5)+(-6*3m)=-78 30-18m=-78 -18m=-78-30 -18m=-108 m=-108/-18 m=6

OpenStudy (anonymous):

-5+3m=-78/-6 =>3m=13+5 =>3m=18 =>m=18/3 =>m=6 okay???

OpenStudy (diyadiya):

@my4babiesandme thats wrong its 30-18m=-78 NOT 30+18m=-78

OpenStudy (anonymous):

m=6

OpenStudy (diyadiya):

its okay :) and its m=6